OpenStudy (gracygirl):

Solve the following equation. Then enter your answer in the space provided using mixed number format. 7(x - 2) = 3(x + 4) This is what I've worked out...

OpenStudy (gracygirl):

This is what I've worked out... 7(x - 2) = 3(x + 4) <--- Original 7x - 14 = 3x + 12 <--- Distrubution 7x - 14 + 14 = 3x + 12 + 14 <--- Move the 14 7x = 3x + 26 7x - 3x = 3x + 26 - 3x <--- Move the 3x 4x = 26 \[\frac{ 4x }{ 4 } = \frac{ 26 }{ 4 }\] x = 6 So why does it want it in mixed number format? And how?

OpenStudy (gracygirl):

@jabez177 @AloneS @Awolflover1

OpenStudy (gracygirl):

The answer (I turned the lesson in) is 6 1/2....but how do you get that from 7(x - 2) = 3(x + 4)? @Will.H

OpenStudy (will.h):

you are wrong! you did everything right However last step when dividing x = 26/4 is equal to 6.5 or 6 1/2 not 6

OpenStudy (will.h):

make sense?

OpenStudy (laura.swart):

7(x - 2) = 3(x + 4) 7x-14=3x+12 4x-14=12 4x=26 x=26/4 which reduces to 13/2

OpenStudy (gracygirl):

Oh, yeah I get it. Thank you both ^_^ :)